The relation 3t=(3x)^1/2 + 6 describes the displacement of a particle in one direction where x is in metres and t is in seconds.The displacement when the velocity is zero is:-

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Hi Divya Thakur

3t=(3x)^1/2 + 6

simplify the expression

x=(1/3)(3t-6)^2.........(1)

for velocity diffrentiate

v=dx/dt=(2/3)(3t-6)*3

                =6t-12

for v=0

  6t-12=0

    or t=2

put value of t in equation 1

x=0

for t<2 motion is not define

motion is just started after t>=2

at this point displacement as well as velocity is zero


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best Divya.
 
Regards,
Askiitians Experts
Badiuddin